Find the equations of the spheres which pass through the circle x²+y²+z²=9, 2x+2y-7=0 and touch the plane x-y+z+3=0
1
Expert's answer
2019-07-16T10:19:49-0400

1. x²+y²+z²=9 is an equation of the sphere S1 with center O1(0,0,0) and radius"R_1=\sqrt{9}=3."


2. The sign distance from the plane ax+by+cz+d=0 to the point X0(x0,y0,z0) is determined by


"D=\frac{ax_0+by_0+cz_0+d}{\sqrt{a^2+b^2+c^2}},"


which is positive if X0 is on the same side of the plane as the normal vector n=(a,b,c) and negative if it is on the opposite side. Hens, the distance from the plane 2x+2y-7=0 to the point O1(0,0,0) is equal to


"D_1=\frac{-7}{\sqrt{2^2+2^2+0^2}}=-\frac{7}{\sqrt{8}}."

3. The radius of the given circle can be determined by the Pythagorean theorem:


"r_1=\sqrt{R_1^2-D_1^2}=\sqrt{3^2-(\frac{-7}{\sqrt{8}})^2}=\sqrt{9-\frac{49}{8}}=\sqrt{\frac{23}{8}}."


4. The sphere S2 passes through the given circle, so its center O2 lies on the line parallel to n1 and passing through O1:


"O_2 = O_1 + λ\bold{n_1}=(0,0,0)+λ(2,2,0)=(2λ,2λ,0),"


and its radius R2 is equal to


"D_{21}=\frac{2*2λ+2*2λ+0*0-7}{\sqrt{2^2+2^2+0^2}}=\frac{8λ-7}{\sqrt{8}};""R_2=\sqrt{r_1^2+D_{21}^2}=\sqrt{\frac{23}{8}+\frac{(8λ-7)^2}{8}}=\sqrt{\frac{23+(8λ-7)^2}{8}}."

5. Because the sphere S2 toches the plane x-y+z+3=0, then R2 is equal to the distance from its center to the plane:


"D_2=\frac{1*2λ-1*2λ+1*0+3}{\sqrt{1^2+(-1)^2+1^2}}=\sqrt{3}=R_2;"

"\sqrt{\frac{23+(8λ-7)^2}{8}}=\sqrt{3};"

"23+(8\u03bb-7)^2=24;"

"(8\u03bb-7)^2=1;"

"8\u03bb-7=\u00b11;"

"8\u03bb=7\u00b11."


6. Equations of the spheres:

1) "λ_1=\frac{6}{8}=\frac{3}{4};" "O_2^1 =(2λ_1,2λ_1,0)=(\frac{3}{2},\frac{3}{2},0);"

"(x-\frac{3}{2})²+(y-\frac{3}{2})²+z²=R_2^2=3;"


2) "λ_2=\frac{8}{8}=1;O_2^2 =(2λ_2,2λ_2,0)=(2,2,0);"

"(x-2)\u00b2+(y-2)\u00b2+z\u00b2=R_2^2=3."


Answer:

"(x-\frac{3}{2})²+(y-\frac{3}{2})²+z²=3;"

"(x-2)\u00b2+(y-2)\u00b2+z\u00b2=3."